OpenStudy (anonymous):

y^6+x^3=y^2+10x ,normal at(0,1) at the given point,find the slope of the curve,the line that is tangent to the curve,or the line that is normal to the curve,as requested.

OpenStudy (phi):

can you be more clear about the question?

OpenStudy (anonymous):

at the given point,find the slope of the curve,the line that is tangent to the curve,or the line that is normal to the curve,as requested.

OpenStudy (anonymous):

\[y ^{6}+x ^{3}=y ^{2}+10x , normal at (0,1)\]

OpenStudy (phi):

I would write the equation as \[ y^6+x^3-y^2-10x=0 \] and use implicit differentiation. solve for dy/dx . can you do that ?

OpenStudy (anonymous):

\[6y ^{5}+3x ^{2}-2y-10=0\]

OpenStudy (phi):

you left out an important part. when you take the derivative of x with respect to x you get \[ \frac{d }{dx}x = \frac{dx}{dx}= 1 \] when you take the derivative of y with respect to x you get \[ \frac{d }{dx}y = \frac{dy}{dx} \] if you have a function of y , for example \(y^2\), you must use the chain rule: \[ \frac{d }{dx}y^2 =2 y \frac{d }{dx}y =2 y \frac{dy }{dx}\]

OpenStudy (phi):

people often abbreviate \( \frac{dy}{dx} \) as y' (easier to type) The derivative is \[ 6y ^{5} y' +3x ^{2}-2y\ y'-10=0 \]

OpenStudy (phi):

now solve for y' (you could solve for y' in terms of x and y , or you could first sub in x=0, y=1 and then solve)